OpenStudy (anonymous):

Write in point-slope form an equation of the line through each pair of points.

OpenStudy (anonymous):

(1,0) and (5,5)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

First you need to find the slope

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

Okay. Now choose one of the plug it in to the point-slope formula

OpenStudy (anonymous):

one of the points*

OpenStudy (anonymous):

can you show me? so i can use this as a base for the rest of the problems

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

thats the formula

OpenStudy (anonymous):

i know the formula

OpenStudy (anonymous):

y-5=5/4(x-5)

OpenStudy (anonymous):

so, once i find the slope, i can choose any of the points i want and just plug them into the formula? and i dont have to use them all?

OpenStudy (anonymous):

Yes, thats correct. No you do not, but only plug numbers into x1/y1 never into x/y

OpenStudy (anonymous):

can i post another one here and you can see if i do it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(-4,10) and (-6,15)

OpenStudy (anonymous):

label x1 y1 and x2 y2...

OpenStudy (anonymous):

slope is 5 over -2?

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

okay, so the formula is y-y1=m(x-x1)

OpenStudy (anonymous):

so would y-10=5/-2(x-15) be an answer?

OpenStudy (anonymous):

No, you have to only use one point. Let's use (-4,10) x= -4 and y=10, so now use this and plug in.

OpenStudy (anonymous):

y-10=5/-2(x-(-4)?

OpenStudy (anonymous):

Remember two negative cancel each other out and it turns into a positive. So in this case it would be x+4

OpenStudy (anonymous):

right. but that would work?

OpenStudy (anonymous):

yes, it would

OpenStudy (anonymous):

thanks bae

OpenStudy (anonymous):

You're welcome :)